# Aesthetic Post-Training Diffusion Models from Generic Preferences with Step-by-step Preference Optimization

### v2

@@ Abstract
Generating visually appealing images is fundamental to modern text-to-image generation models. A potential solution to better aesthetics is direct preference optimization (DPO), which has been applied to diffusion models to improve general image quality including prompt alignment and aesthetics. Popular DPO methods propagate preference labels from clean image pairs to all the intermediate steps along the two generation trajectories. However, preference labels provided in existing datasets are blended with layout and aesthetic opinions, which would disagree with aesthetic preference. Even if aesthetic labels were provided (at substantial cost), it would be hard for the two-trajectory methods to capture nuanced visual differences at different steps. To improve aesthetics economically, this paper uses existing generic preference data and introduces step-by-step preference optimization (SPO) that discards the propagation strategy and allows fine-grained image details to be assessed. Specifically, at each denoising step, we 1) sample a pool of candidates by denoising from a shared noise latent, 2) use a step-aware preference model to find a suitable win-lose pair to supervise the diffusion model, and 3) randomly select one from the pool to initialize the next denoising step. This strategy ensures that the diffusion models to focus on the subtle, fine-grained visual differences instead of layout aspect. We find that aesthetic can be significantly enhanced by accumulating these improved minor differences. When fine-tuning Stable Diffusion v1.5 and SDXL, SPO yields significant improvements in aesthetics compared with existing DPO methods while not sacrificing image-text alignment compared with vanilla models. Moreover, SPO converges much faster than DPO methods due to the step-by-step alignment of fine-grained visual details. Code and models are available at https://github.com/RockeyCoss/SPO.
Figure 1: Problem illustration of existing DPO methods. We show two denoising trajectories of a preferred image (a) and a dispreferred image (b) generated from prompt “A cat jumps on a dog”. (1) Disagreement between general preference and aesthetic preference. While (a) is preferred due to its better layout, its details are poorer than (b). Red boxes show the erroneous fusion of the cat’s leg and the dog’s body. (2) Large sample differences between trajectories. This makes it difficult to identify subtle aesthetics related differences at each step.
@@ 1 Introduction
However, for two problems existing systems are not best effective for aesthetic alignment. First, general preferences provided in public datasets like Pick-a-Pic [11] are not aesthetics specific, and they often disagree. In Fig. 1, (a) is generally preferred because of the correctly generated dog, cat, and their spatial arrangement, but in terms of aesthetics, (a) should be dispreferred. More examples are provided in the supplementary material. This introduces noisy supervision signal and compromises the model improvement towards better aesthetics. Technically, this problem can be fixed by manually annotating an aesthetics-only preference dataset, but it is very expensive, because of layout influence and complexity of aesthetics. Second, in the two-trajectory methods, image pairs at each denoising step look very different, as shown in Fig. 1. Even if accurate aesthetic preferences were provided at each step (at great cost), it would still be very non-trivial to learn, because the large layout discrepancy would dominate over aesthetic nuances.
To better align diffusion models with aesthetic preference, we introduce step-by-step preference optimization (SPO). SPO is new in that it completely discards the current preference propagation strategy, pushing for evaluating image details. Specifically, at each step beginning with a noisy image $\boldsymbol{x}_{t}$, we generate a pool of $\boldsymbol{x}_{t-1}$ samples. We evaluate their quality using a step-aware preference model (SPM) and assign win/lose labels to the pair showing the largest quality difference. We then randomly select an image from the pool to initialize timestep $t-1$. Because 1) the win-lose pair comes from the same image and 2) is generated in one or very few steps, the two samples would only exhibit small differences in details. SPM allows us to capture such details differences and guide the diffusion model to generate more visually pleasing images.
We use SPO to fine-tune Stable Diffusion v1.5 (SD-1.5) [23] and SDXL [18]. Although we use a training set with generic human perferences [11], we demonstrate significant improvement in aesthetics compared with those fine-tuned by popular DPO methods. Moreover, SPO is much faster to converge compared with DPO. This is because the step-by-step design makes it easier to focus on fine-grained visual details. We summarize key points of this paper below.
@@
  • We point out that generic human preferences are often different from pure aesthetic preferences and that obtaining aesthetic-only preference data is very expensive.
  • We design SPO, where we determine win-lose pairs at each step in an online manner. We make sure win-lose pairs come from the same noisy sample, so after one or very few steps of denoising their difference is small and in fine-grained details, and can be captured by a preference model trained with generic preference data.
@@ 2 Related Work
Recently, inspired by post-training methods that improve LLMs, e.g., reinforcement learning from human feedback (RLHF) [16, 35], various post-training methods are proposed to align pre-trained diffusion models with human preferences. For example, Chen et al. [2] leverage the PPO loss [26] to fine-tune the text encoder of diffusion models. AligningT2I [13] develops a reward model to evaluate the quality of generated images and fine-tunes the diffusion model using image-text pairs, weighted by assessment of the reward model. DPOK [7] and DDPO [1] use policy gradient to fine-tune diffusion models, aiming at maximizing reward signals. Furthermore, ReFL [31], DRaFT [3], and AlignProp [19] directly propagate the gradients through differentiable reward models to fine-tune the denoising steps. Recent methods are inspired by direct preference optimization (DPO) [21], which post-trains LLMs by eliminating the need for explicit reward models. Diffusion-DPO [29] fine-tunes diffusion models on the Pick-a-Pic [11] dataset that contains image preference pairs. D3PO [32] generates pairs of images from the same prompt and uses a preference model to identify preferred and dispreferred images. DenseReward [33] improves the DPO scheme with a temporal discounting approach to emphasize initial denoising steps. These methods optimize the trajectory-level preference, where the accumulated image differences are too large to allow the network to focus on aesthetics subtleties. In comparison, SPO by its step-by-step mechanism can focus on nuanced visual differences in just a single or few steps.
@@ 3 DPO Revisit and Diagnosis
General formulation. Given a generation model $\pi_{\theta}(\cdot)$ and condition $c$, the probability of generating output $o$ is $\pi_{\theta}(o\mid c)$. We use $\pi_{\theta}(\cdot)$ to generate a set of output pairs $\mathcal{S}$, where each pair comes from the same condition $c$. Human or preference model is employed to label the preference order of the output pairs as $(o^{w},o^{l},c)$, where $o^{w}$ is the preferred output and $o^{l}$ is dispreferred. According to Rafailov et al. [21], the DPO loss used to fine-tune $\pi_{\theta}$ is defined as:
@@
$\pi_{\textrm{ref}}(\cdot)$ refers to the reference model, which is usually pre-trained. $\sigma(\cdot)$ is the sigmoid function. $\beta$ is a scalar controlling the strength of regularization.
DPO for diffusion model post-training. We denote text-to-image diffusion models with parameters $\theta$ as $p_{\theta}$ and text prompt as $c$. The denoising process generates intermediate states $\{\boldsymbol{x}_{T},\boldsymbol{x}_{T-1},...,\boldsymbol{x}_{1},\boldsymbol{x}_{0}\}$. Existing works including Diffusion-DPO [29] and D3PO [32] measure preference based on the final image $\boldsymbol{x}_{0}$ and assign the preference for $\boldsymbol{x}_{0}$ directly to the entire generation trajectory, or all the intermediate states. Let $\mathcal{T}_{w}$ and $\mathcal{T}_{l}$ denote the denoising trajectories which generate the preferred and dispreferred images, respectively. Using the Markov property of diffusion models and Jensen’s inequality, they reformulate the general DPO loss in Eq. 1 into the following step-wise form:
@@
Figure 3: Comparing frameworks of SPO and the existing Diffusion-DPO and D3PO approaches. SPO does not adopt direct preference propagation like the DPO methods. In SPO, a pool of samples are generated at each step, from which a proper win/lose pair is selected and used to fine-tune the diffusion model. Then, a single sample is randomly selected to initialize the next iteration.
Diagnosis on aesthetic alignment. Aligning diffusion models with human aesthetic preference is very challenging. From architecture, win-lose pairs in the two-trajectory methods usually differ significantly (primarily in layout), rendering it hard to focus on image detail comparisons. From data, existing aesthetic scoring datasets [15, 25] do not provide paired image data coming from the same prompt. Collecting a dedicated aesthetic preference dataset would be more costly compared with tasks like classification or prompt alignment preference. So a cost-efficient choice is existing generic preference datasets where annotators were asked score images based on holistic opinions like prompt alignment and aesthetics. However as pointed out in Section 1, generic preferences may be contradictory to aesthetic preference. If we directly propagate them to all diffusion steps like Diffusion-DPO and D3PO do (see Fig. 3), noisy preferences will compromise fine-tuning.
@@ 4.1 Framework Overview
To align diffusion models with aesthetic preference while still using generic preferences, we propose step-by-step preference optimization (SPO), an online reinforcement learning method. Fig. 3 $(c)$ depicts its workflow. Given an intermediate $\boldsymbol{x}_{t}$, at timestep $(t{-}1)$, we sample a pool of denoised samples $\{\boldsymbol{x}_{t-1}^{1},\boldsymbol{x}_{t-1}^{2},\cdots,\boldsymbol{x}_{t-1}^{k}\}$. We apply a step-aware preference model (Section 4.2) to compare the quality of these candidate samples, and select the highest-quality sample and the lowest-quality sample as the win-sample and lose-sample, respectively. SPO ensures the win-lose pair comes from the same $\boldsymbol{x}_{t}$ and thus have small differences in image details to reflect aesthetics. We then randomly select a sample from the candidate pool (Section 4.3), which is used to initialize the next iteration. SPO is optimized by a revised DPO loss function (Section 4.4) and can be extended to the SDXL model (Section 4.5).
@@ 4.2 Step-Aware Preference Model (SPM)
For SPM training, we use the same objective as [4], making models useful for noisy images. Specifically, we add the same amount of noise to a pair of clean images, assuming that their preference order can be kept. During training, for each pair of images and their human-labeled preference, we randomly sample a timestep $t$ and add the same noise to both images to obtain $\boldsymbol{x}^{w}_{t}$ and $\boldsymbol{x}^{l}_{t}$. Then we feed the noisy intermediate pair $\{\boldsymbol{x}^{w}_{t},\boldsymbol{x}^{l}_{t}\}$ and timestep $t$ to SPM and train the model to correctly predict the preference, using loss function $\mathcal{L}_{\text{pref}}=\left(\log 1-\log\hat{p}_{w}\right)$, where $\hat{p}_{w}$ is the probability of the win image being the preferred one, following [11]. $\hat{p}_{w}$ is computed using the following equation:
@@ 4.3 Random Selection of $\boldsymbol{x}_{t}$ to Start Next Step
In SPO, we randomly sample $\boldsymbol{x}_{t}$ from the candidate pool $\{\boldsymbol{x}_{t}^{1},\boldsymbol{x}_{t}^{2},\cdots,\boldsymbol{x}_{t}^{k}\}$, which is then used to start the next denoising step. This ensures every win-lose pair comes from the same latent. Since $\boldsymbol{x}_{t}$ is very noisy when $t$ is large, we use the same $\boldsymbol{x}_{t}$ when $t$ is greater than a threshold $\kappa$. Only when $t\leq\kappa$ do we apply random selection. This random selection is shown in Fig. 3 $(c)$: selection of $\boldsymbol{x}_{t-1}$ is depicted and is consistent with the above discussion.
@@ 4.4 Objective Function of SPO
where $\mathcal{U}$ and $\mathcal{N}$ are the uniform distribution and Gaussian distributions, respectively.
@@ 4.5 Extension to Multi-step Preference Optimization for SDXL
For stronger models like SDXL, we observe that the difference between $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t-1}$ is too small, so is the difference between the selected candidates $\boldsymbol{x}^{w}_{t-1}$ and $\boldsymbol{x}^{l}_{t-1}$. While a relatively small difference allows SPM to focus on image details, that difference is too small would create ambiguities and confuse fine-tuning. To address this issue, we expand step-by-step preference optimization to multi-step preference optimization (MSPO), which uses multiple denoising processes to increase the diversity of the candidate set (see Fig. 4). This simple extension allows us to select more different samples and more clear preference signals.
@@ 4.6 Discussions and Insights
How often does generic preference and aesthetics preference disagree? It is hard to quantify because it is non-trivial to annotate a validation set. Fig. 1 and supplementary material depict some disagreement scenarios based on careful manual inspection. In fact, because of the nuanced nature of aesthetics, an image may be superior in certain aesthetic aspects while its paired image are better in other aspects. So there are probably more disagreement scenarios than we can think of now.
@@ 5.1 Experimental Setup
Evaluation Protocol. This paper evaluates image quality with four automatic metrics. We use PickScore [11], HPSV2 [30] and ImageReward [31] as prompt-aware human preference estimate. These models are trained on human preference datasets and learn to replicate human decisions about which images are more favorable. We use Aesthetic score [25] to evaluate visual appeal. This score is prompt agnostic and employs a linear estimator on top of the vision encoder of CLIP [20] to predict the aesthetic quality of images. Note that PickScore, HPSV2, and ImageReward all assess aesthetics to some extent. Apart from these automatic metrics, we also use human studies. We invite 10 people to assess 300 pairs of images generated by two methods of interest. Their preference is summarized into winning percentage from 0 to 100%. The text prompts are randomly selected from PartiPrompts (100 prompts) [34] and HPSV2 benchmark (200 prompts) [30].
@@
Table 2: Comparing AI feedback methods on SD-1.5.
@@
Table 3: Method comparison on GenBench based on SDXL. $\dagger$: results are reproduced with a classifier-free guidance scale of 5.0 and 50 inference steps. GenBench evaluates image-text alignment.
@@
Implementation details. We apply DDIM [28] with $\eta=1.0$ and 20 timesteps as the sampler and use classifier free guidance [9] with scale 5.0 for sampling during online training. We use LoRA [10] for both SD-1.5 and SDXL, fine-tuning the models for $10$ epochs. The LoRA rank is 4 and 64 for SD-1.5 and SDXL, respectively. We set the strength of regularization $\beta=10$. For SD-1.5, we set the batch size as $40$ and learning rate as $6e^{-5}$. For SDXL, we set the batch size as $8$, gradient accumulation as $2$, and learning rate as $1e^{-5}$. Since very noisy images are difficult to compare, we do not use SPM to very early stages. That is, we only compute the preference of $\boldsymbol{x}_{t}$ when $t\leq\kappa$ and consider all $\boldsymbol{x}_{t}$ with $t>\kappa$ as tied. We empirically set $\kappa$ as $750$ and will evaluate $\kappa$ in Section 5.4. When fine-tuning SDXL with the MSPO (Section 4.5), we set the number of inner steps to $4$. We do not apply SDXL refiner [18] to ensure fair comparison. For SPM training, we adopt learning rates of 3e-6 and 1e-6 for SD-1.5 and SDXL, respectively. We use DDIM scheduler with classifier free guidance scale of 5 and 20 steps to inference on validation prompts.
@@
Table 7: Impact of #inner steps when fine-tuning SDXL with MSPO. We set $j=4$.

### v3

@@ Abstract
Generating visually appealing images is fundamental to modern text-to-image generation models. A potential solution to better aesthetics is direct preference optimization (DPO), which has been applied to diffusion models to improve general image quality including prompt alignment and aesthetics. Popular DPO methods propagate preference labels from clean image pairs to all the intermediate steps along the two generation trajectories. However, preference labels provided in existing datasets are blended with layout and aesthetic opinions, which would disagree with aesthetic preference. Even if aesthetic labels were provided (at substantial cost), it would be hard for the two-trajectory methods to capture nuanced visual differences at different steps. To improve aesthetics economically, this paper uses existing generic preference data and introduces step-by-step preference optimization (SPO) that discards the propagation strategy and allows fine-grained image details to be assessed. Specifically, at each denoising step, we 1) sample a pool of candidates by denoising from a shared noise latent, 2) use a step-aware preference model to find a suitable win-lose pair to supervise the diffusion model, and 3) randomly select one from the pool to initialize the next denoising step. This strategy ensures that diffusion models focus on the subtle, fine-grained visual differences instead of layout aspect. We find that aesthetics can be significantly enhanced by accumulating these improved minor differences. When fine-tuning Stable Diffusion v1.5 and SDXL, SPO yields significant improvements in aesthetics compared with existing DPO methods while not sacrificing image-text alignment compared with vanilla models. Moreover, SPO converges much faster than DPO methods due to the use of more correct preference labels provided by the step-aware preference model. Code and models are available at https://github.com/RockeyCoss/SPO.
Figure 1: Problem illustration of existing DPO methods. We show two denoising trajectories of a preferred image (a) and a dispreferred image (b) generated from prompt “A cat jumps on a dog”. (1) Disagreement between generic preferences and aesthetic preference. While (a) is preferred due to its better layout, its details are poorer than (b). Red boxes show the erroneous fusion of the cat’s leg and the dog’s body. (2) Large sample differences between trajectories. This makes it difficult to identify subtle aesthetic-related differences at each step.
@@ 1 Introduction
However, for two problems existing systems are not best effective for aesthetic alignment. First, generic preferences provided in public datasets like Pick-a-Pic [11] are not aesthetics specific, and they often disagree. In Fig. 1, (a) is generally preferred because of the correctly generated dog, cat, and their spatial arrangement, but in terms of aesthetics, (a) should be dispreferred. More examples are provided in the supplementary material. This introduces noisy supervision signals and compromises the model improvement towards better aesthetics. Technically, this problem can be fixed by manually annotating an aesthetics-only preference dataset, but it is very expensive, because of layout influence and complexity of aesthetics. Second, in two-trajectory methods, the images within each pair at each denoising step look very different, as shown in Fig. 1. Even if accurate aesthetic preference was provided at each step (at great cost), it would still be very non-trivial to learn, because the large layout discrepancy would dominate over aesthetic nuances.
To better align diffusion models with aesthetic preference, we introduce step-by-step preference optimization (SPO). SPO is new in that it completely discards the current preference propagation strategy, pushing for evaluating image details. Specifically, at each step beginning with a noisy image $\boldsymbol{x}_{t}$, we generate a pool of $\boldsymbol{x}_{t-1}$ samples. We evaluate their quality using a step-aware preference model (SPM) and assign win/lose labels to the pair showing the largest quality difference. We then randomly select an image from the pool to initialize timestep $t-1$. Because the win-lose pair 1) comes from the same image and 2) is generated in one or very few steps, the two samples would only exhibit small differences in details. SPM allows us to capture such detailed differences and guide the diffusion model to generate more visually pleasing images.
We use SPO to fine-tune Stable Diffusion v1.5 (SD-1.5) [23] and SDXL [18]. Although we use a training set with generic human perferences [11], we demonstrate significant improvement in aesthetics compared with those fine-tuned by popular DPO methods. Moreover, SPO converges much faster than Diffusion-DPO. This is because the step-by-step design makes it easier to focus on fine-grained visual details, and SPM produces more accurate preference labels. We summarize key points of this paper below.
@@
  • We point out that generic human preferences are often different from pure aesthetic preference and that obtaining aesthetic-only preference data is very expensive.
  • We design SPO, where we determine win-lose pairs at each step in an online manner. We make sure win-lose pairs come from the same noisy sample, so after one or very few steps of denoising their differences are small and in fine-grained details, and can be captured by a preference model trained with generic preference data.
@@ 2 Related Work
Recently, inspired by post-training methods that improve LLMs, e.g., reinforcement learning from human feedback (RLHF) [16, 35], various post-training methods are proposed to align pre-trained diffusion models with human preferences. For example, Chen et al. [2] leverage the PPO loss [26] to fine-tune the text encoder of diffusion models. AligningT2I [13] develops a reward model to evaluate the quality of generated images and fine-tunes the diffusion model using image-text pairs, weighted by assessment of the reward model. DPOK [7] and DDPO [1] use policy gradient to fine-tune diffusion models, aiming at maximizing reward signals. Furthermore, ReFL [31], DRaFT [3], and AlignProp [19] directly propagate the gradients through differentiable reward models to fine-tune the denoising steps. Recent methods are inspired by direct preference optimization (DPO) [21], which eliminates the need for explicit reward models when post-training LLMs. Diffusion-DPO [29] fine-tunes diffusion models on the Pick-a-Pic [11] dataset containing image preference pairs. D3PO [32] generates pairs of images from the same prompt and uses a preference model to identify preferred and dispreferred images. DenseReward [33] improves the DPO scheme with a temporal discounting approach to emphasize initial denoising steps. These methods optimize the trajectory-level preference, where the accumulated image differences are too large to allow the network to focus on aesthetics subtleties. In comparison, SPO by its step-by-step mechanism can focus on nuanced visual differences in just a single or few steps.
@@ 3 DPO Revisit and Diagnosis
General formulation. Given a generation model $\pi_{\theta}(\cdot)$ and a condition $c$, the probability of generating output $o$ is $\pi_{\theta}(o\mid c)$. We use $\pi_{\theta}(\cdot)$ to generate a set of output pairs $\mathcal{S}$, where each pair comes from the same condition $c$. Human or preference model is employed to label the preference order of the output pairs as $(o^{w},o^{l},c)$, where $o^{w}$ is the preferred output and $o^{l}$ is dispreferred. According to Rafailov et al. [21], the DPO loss used to fine-tune $\pi_{\theta}$ is defined as:
@@
$\pi_{\textrm{ref}}(\cdot)$ and $\sigma(\cdot)$ refer to the reference model and the sigmoid function, respectively. $\beta$ is the strength of regularization.
DPO for diffusion model post-training. We denote the text-to-image diffusion model with parameters $\theta$ as $p_{\theta}$ and text prompt as $c$. The denoising process generates intermediate states $\{\boldsymbol{x}_{T},\boldsymbol{x}_{T-1},...,\boldsymbol{x}_{1},\boldsymbol{x}_{0}\}$. Existing works including Diffusion-DPO [29] and D3PO [32] measure preference based on the final image $\boldsymbol{x}_{0}$ and assign the preference for $\boldsymbol{x}_{0}$ directly to the entire generation trajectory, or all the intermediate states. Let $\mathcal{T}_{w}$ and $\mathcal{T}_{l}$ denote the denoising trajectories which generate the preferred and dispreferred images, respectively. Using the Markov property of diffusion models and Jensen’s inequality, they reformulate the general DPO loss in Eq. 1 into the following step-wise form:
@@
Figure 3: Comparing frameworks of SPO, Diffusion-DPO, and D3PO approaches. SPO does not adopt direct preference propagation as other DPO methods do. In SPO, a pool of samples are generated at each step, from which a proper win/lose pair is selected and used to fine-tune the diffusion model. Then, a single sample is randomly selected to initialize the next iteration.
Diagnosis on aesthetic alignment. Aligning diffusion models with human aesthetic preference is very challenging. From architecture, win-lose pairs in the two-trajectory methods usually differ significantly (primarily in layout), rendering it hard to focus on image detail comparisons. From data, existing aesthetic scoring datasets [15, 25] do not provide paired image data coming from the same prompt. Collecting a dedicated aesthetic preference dataset would be more costly compared with tasks like classification or prompt alignment preference. So a cost-efficient choice is existing generic preference datasets where annotators were asked to score images based on holistic opinions like prompt alignment and aesthetics. However as pointed out in Section 1, generic preferences may be contradictory to aesthetic preference. If we directly propagate them to all diffusion steps like Diffusion-DPO and D3PO do (see Fig. 3), noisy preferences will compromise fine-tuning.
@@ 4.1 Framework Overview
To align diffusion models with aesthetic preference while still using generic preferences, we propose step-by-step preference optimization (SPO), an online reinforcement learning method. Fig. 3 $(c)$ depicts its workflow. Given an intermediate $\boldsymbol{x}_{t}$, at timestep $t$, we sample a pool of denoised samples $\{\boldsymbol{x}_{t-1}^{1},\boldsymbol{x}_{t-1}^{2},\cdots,\boldsymbol{x}_{t-1}^{k}\}$. We apply a step-aware preference model (Section 4.2) to compare the quality of these candidate samples, and select the highest-quality sample and the lowest-quality sample as the win-sample and the lose-sample, respectively. SPO ensures the win-lose pair comes from the same $\boldsymbol{x}_{t}$ and thus have small differences in image details to reflect aesthetics. We then randomly select a sample from the candidate pool (Section 4.3), which is used to initialize the next iteration. SPO is optimized by a revised DPO loss function (Section 4.4) and can be extended to the SDXL model (Section 4.5).
@@ 4.2 Step-Aware Preference Model (SPM)
For SPM training, we initialize the model with PickScore [11] and fine-tune it following [4], making the model useful for noisy images. Specifically, we add the same amount of noise to a pair of clean images, assuming that their preference order can be kept. During training, for each pair of images and their human-labeled preference, we randomly sample a timestep $t$ and add the same noise to both images to obtain $\boldsymbol{x}^{w}_{t}$ and $\boldsymbol{x}^{l}_{t}$. Then we feed the noisy intermediate pair $\{\boldsymbol{x}^{w}_{t},\boldsymbol{x}^{l}_{t}\}$ and timestep $t$ to SPM and train the model to correctly predict the preference, using loss function $\mathcal{L}_{\text{pref}}=\left(\log 1-\log\hat{p}_{w}\right)$, where $\hat{p}_{w}$ is the probability of the win image being the preferred one, following [11]. $\hat{p}_{w}$ is computed using the following equation:
@@ 4.3 Random Selection of $\boldsymbol{x}_{t}$ to Start Next Step
In SPO, we randomly sample $\boldsymbol{x}_{t}$ from the candidate pool $\{\boldsymbol{x}_{t}^{1},\boldsymbol{x}_{t}^{2},\cdots,\boldsymbol{x}_{t}^{k}\}$, which is then used to start the next denoising step. This ensures every win-lose pair comes from the same latent. Since $\boldsymbol{x}_{t}$ is very noisy when $t$ is large, we use the standard diffusion sampling process when $t$ is greater than the threshold $\kappa$. Only when $t\leq\kappa$ do we sample candidate pools and apply random selection. This random selection is shown in Fig. 3 $(c)$: selection of $\boldsymbol{x}_{t-1}$ is depicted and is consistent with the above discussion.
@@ 4.4 Objective Function of SPO
where $\mathcal{U}$ and $\mathcal{N}$ are the uniform distribution and Gaussian distribution, respectively.
@@ 4.5 Extension to Multi-step Preference Optimization for SDXL
For stronger models like SDXL, we observe that the difference between $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t-1}$ is too small, so is the difference between the selected candidates $\boldsymbol{x}^{w}_{t-1}$ and $\boldsymbol{x}^{l}_{t-1}$. While a relatively small difference allows SPM to focus on image details, that difference is too small would create ambiguities and confuse fine-tuning. To address this issue, we expand step-by-step preference optimization to multi-step preference optimization (MSPO), which uses multiple denoising steps to increase the diversity of the candidate set (see Fig. 4). This simple extension allows us to select more different samples and more clear preference signals.
@@ 4.6 Discussions and Insights
How often does generic preferences and aesthetic preference disagree? It is hard to quantify because it is non-trivial to annotate a validation set. Fig. 1 and supplementary material depict some disagreement scenarios based on careful manual inspection. In fact, because of the nuanced nature of aesthetics, an image may be superior in certain aesthetic aspects while its paired image is better in other aspects. So there are probably more disagreement scenarios than we can think of now.
@@ 5.1 Experimental Setup
Evaluation Protocol. This paper evaluates image quality with four automatic metrics. We use PickScore [11], HPSV2 [30] and ImageReward [31] for prompt-aware human preference estimation. These models are trained on human preference datasets and learn to replicate human decisions about which images are more favorable. We use Aesthetic score [25] to evaluate visual appeal. This score is prompt agnostic and employs a linear estimator on top of the vision encoder of CLIP [20] to predict the aesthetic quality of images. Note that PickScore, HPSV2, and ImageReward all assess aesthetics to some extent. Apart from these automatic metrics, we also use human studies. We invite 10 people to assess 300 pairs of images generated by two methods of interest. Their preferences are summarized into winning percentage from 0 to 100%. The text prompts are randomly selected from PartiPrompts (100 prompts) [34] and HPSV2 benchmark (200 prompts) [30].
@@
Table 2: Comparing different methods on SD-1.5.
@@
Table 3: Method comparison on GenEval based on SDXL. $\dagger$: results are reproduced with a classifier-free guidance scale of 5.0 and 50 inference steps. GenEval evaluates image-text alignment.
@@
Implementation details. We apply DDIM [28] with $\eta=1.0$ and 20 timesteps as the sampler and use classifier free guidance [9] with scale 5.0 for sampling during online training. We use LoRA [10] for both SD-1.5 and SDXL, fine-tuning the models for $10$ epochs. The LoRA rank is 4 and 64 for SD-1.5 and SDXL, respectively. We set the strength of regularization $\beta=10$. For SD-1.5, we set the batch size as $40$ and learning rate as $6e^{-5}$. For SDXL, we set the batch size as $8$, gradient accumulation as $2$, and learning rate as $1e^{-5}$. Since very noisy images are difficult to compare, we do not use SPM to very early stages. That is, we only compute the preference of $\boldsymbol{x}_{t}$ when $t\leq\kappa$ and consider all $\boldsymbol{x}_{t}$ with $t>\kappa$ as tied. We empirically set $\kappa$ as $750$ and will evaluate $\kappa$ in Section 5.4. When fine-tuning SDXL with the MSPO (Section 4.5), we set the number of inner steps to $4$. We do not apply SDXL refiner [18] to ensure fair comparison. For SPM training, we adopt learning rates of 3e-6 and 1e-6 for SD-1.5 and SDXL, respectively. We use DDIM scheduler with classifier free guidance scale of 5 and 20 steps to perform inference on validation prompts.
@@
Table 7: Impact of #inner steps $j$ when fine-tuning SDXL with MSPO. We set $j=4$.